OpenStudy (anonymous):

Simplify the square root of 5x-the the square root of 20x- the square root of 80x

OpenStudy (anonymous):

\[\sqrt{5x}+2\sqrt{5x}-4\sqrt{5x}\]

OpenStudy (anonymous):

\[\sqrt{5x}-\sqrt{20x}-\sqrt{80x}\]

OpenStudy (anonymous):

combine like terms go get \[-\sqrt{5x}\]